OpenStudy (anonymous):

multiply \[2/3a^{4}(6a ^{5}-12a ^{3}-5/8)\]

OpenStudy (unklerhaukus):

\[\frac{2}{3a^4} (6a^5−12a^3−\frac{5}{8})\]\[=4a−\frac{8}{a}−\frac{10}{24a^4}\]

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

\[-\frac{5}{12a^4}-\frac{8}{a}+4a\]